### v1

@@ E.1 Varying sampling methods
Fig 18, Fig 19, and Fig 20 report creativity, memorization and coherence (i.e., fraction of generated strings that are coherent) for various sampling methods (greedy decoding and nucleus sampling) with various prefix conditionings (namely, null, pause and hash).
Figure 18: Algorithmic creativity under various sampling conditions for Gemma v1 (2B): Across all conditions, and in almost all datasets (with a few exceptions in Line Construction), multi-token prediction improves creativity. Furthermore, hash-conditioning achieves best algorithmic creativity, with a longer hash helping more.
@@ F.1 Ablation studies
F.2 Effect of hash string length
We provide an ablation study on the hash string length for NTP vs MTP on the Sibling Discovery task (Fig 23). We see that longer hash strings lead to higher algorithmic creativity.
Figure 23: GPT-2 (86M) Transformer achieves higher algorithmic creativity with longer hash strings. We report algorithmic creativity with hash strings of length 4 and 10, with both NTP and teacherless MTP.
F.3 Format sensitivity for Triangle Discovery
@@ H.2 Decomposing algorithmic creativity as diversity and memorization
Better creativity can be achieved either by enhancing diversity or by reducing memorization – we try to disentangle these factors in this section. In Fig 27, we plot the algorithmic creativity, diversity, and memorization scores at the checkpoint of best algorithmic creativity. We see that hash-conditioning contributes to higher diversity but does little to bring down memorization; however, teacherless training contributes to higher diversity and also to reducing memorization. In Fig 26, we see that the best creativity and best diversity are not achieved at the same checkpoint.
@@
Figure 27: Decomposition of algorithmic creativity for GPT-2 (86M) in Sibling Discovery: We report algorithmic creativity, diversity and memorization at the checkpoint of best algorithmic creativity. We see that hash-conditioning contributes to higher diversity but does not help bring down memorization; teacherless training helps both diversity and in bringing down memorization.
@@ I.1 Additional graphs for effect of multi-token training
I.2 Effect of hash-conditioning
We also conducted hash-conditioning experiments as described in §3.1. The hash strings we use are $10$ randomly sampled uppercase characters from the English alphabet. We report the quality-diversity plots in Fig 31 (for next-token prediction on XSUM) and Fig 32 (for multi-token prediction on XSUM). As such, we do not find any changes in diversity, perhaps because this is not a sufficiently open-ended task.
@@ I.2 Effect of hash-conditioning
Figure 31: Hash-conditioning has no effect on diversity for GPT models on XSUM summarization with next-token prediction.
@@
Figure 32: Hash-conditioning has no effect on diversity for GPT models on XSUM summarization with multi-token prediction.
@@
Other studies have proposed benchmarks for evaluating creativity. AidanBench McLaughlin et al. (2024) and NoveltyBench Zhang et al. (2025) evaluate LMs on their ability to produce diverse and coherent responses by penalizing repetition across generations. However, they do not measure originality relative to training data, leaving open whether outputs are genuinely novel or simply unseen paraphrases/recombinations. Zhao et al. (2024) evaluate LM creativity using the Torrance Tests of Creative Thinking, a standard in human psychometrics. Another line of work such as Alchemy Wang et al. (2021), IVRE Xu et al. (2022), and DiscoveryWorld Jansen et al. (2024) present simulations with hidden facts and rules, requiring LMs to explore, hypothesize, and test through interaction. While these simulations focus on pretrained models rather than examining how training shapes creative capabilities, they serve as valuable and realistic benchmarks for assessing the role of creativity in scientific discovery.
@@
In support of next-token prediction, there are arguments (Shannon, 1948; 1951; Alabdulmohsin et al., 2024) that claim that language is captured by NTP with models even superceding humans (Shlegeris et al., 2022) at NTP. There are also theoretical results emphasizing the expressivity (Merrill & Sabharwal, 2024; Feng et al., 2023) and learnability (Malach, 2023; Wies et al., 2023) of autoregressive Transformers as long as there is a sufficiently long chain of thought.
@@
While these methods employ diverse strategies, a common feature is their reliance on multi-token objectives that capture broader dependencies across entire sequences. Representative examples include teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) and independent output heads or modules (Gloeckle et al., 2024; DeepSeek-AI et al., 2024) or inserting a lookahead attention (Du et al., 2023). Another line of research is discrete diffusion models (Hoogeboom et al., 2021; Austin et al., 2021; Gong et al., 2023; Lou et al., 2023), which avoid strict left-to-right factorization by iteratively refining an entire sequence at multiple positions. There are other models as well, such as energy-based models Dawid & LeCun (2023) and non-autoregressive models or (Gu et al., 2018).
@@
Graph tasks have been used to understand various limitations of Transformers in orthogonal settings. Bachmann & Nagarajan (2024); Saparov et al. (2024) report that Transformers are limited in terms of learning to search tasks on graphs, while Sanford et al. (2024) provide positive expressivity results for a range of algorithmic tasks that process an graph. These works differ from our study of combinational creativity since their graphs are provided in-context and the tasks have a unique answer. Other works (Schnitzler et al., ; Yang et al., 2024a; b) study multi-hop question answering on a knowledge graph; however, this does not require planning.
@@
One line of work relevant to us in the history of generative models is RNN-based VAE for text data (Bowman et al., 2016). The motivation, like in our work, was to learn high-level semantic features rather than next-token features with the hope of producing more novel sentences. However, this suffered from posterior collapse, where the model ignores the latent variable altogether inspiring various solutions (Yang et al., 2017; Goyal et al., 2017). Our results on hash-conditioning are also reminiscent of a line of work on exploration in reinforcement learning (RL), where it has been shown that adding noises to the policy model parameters enables more efficient exploration than directly adding noises to the output space Plappert et al. (2017); Fortunato et al. (2017).

### v2

@@ E.1 Varying sampling methods
Fig 18, Fig 19, and Fig 20 report creativity, memorization and coherence (i.e., fraction of generated strings that are coherent) for various sampling methods (greedy decoding and nucleus sampling) with various prefix conditionings (namely, null, pause and seed).
Figure 18: Algorithmic creativity under various sampling conditions for Gemma v1 (2B): Across all conditions, and in almost all datasets (with a few exceptions in Line Construction), multi-token prediction improves creativity. Furthermore, seed-conditioning achieves best algorithmic creativity, with a longer seed helping more.
@@ F.1 Ablation studies
F.2 Effect of seed length
We provide an ablation study on the seed length for NTP vs MTP on the Sibling Discovery task (Fig 23 (left)). We see that longer seeds lead to higher algorithmic creativity.
Refer to caption    Refer to caption
Figure 23: GPT-2 (86M) Transformer achieves higher algorithmic creativity with longer seeds (top): We report algorithmic creativity with seeds of length 4 and 10, with both NTP and teacherless MTP. we do not see gains of seed-conditioning when it comes to diffusion training (right): We report algorithmic creativity of the SEDD (90M) model with or without seed-conditioning.
F.3 Seed-conditioning for diffusion training
In Fig 23 (right), we apply seed-conditioning to the SEDD (90M) on the Sibling Discovery task. Unlike what we see on transformers, we do not observe improved creativity when applying seed-conditioning to diffusion training.
F.4 Format sensitivity for Triangle Discovery
@@ H.2 Decomposing algorithmic creativity as diversity and memorization
Better creativity can be achieved either by enhancing diversity or by reducing memorization – we try to disentangle these factors in this section. In Fig 27, we plot the algorithmic creativity, diversity, and memorization scores at the checkpoint of best algorithmic creativity. We see that seed-conditioning contributes to higher diversity but does little to bring down memorization; however, teacherless training contributes to higher diversity and also to reducing memorization. In Fig 26, we see that the best creativity and best diversity are not achieved at the same checkpoint.
@@
Figure 27: Decomposition of algorithmic creativity for GPT-2 (86M) in Sibling Discovery: We report algorithmic creativity, diversity and memorization at the checkpoint of best algorithmic creativity. We see that seed-conditioning contributes to higher diversity but does not help bring down memorization; teacherless training helps both diversity and in bringing down memorization.
@@ I.1 Additional graphs for effect of multi-token training
I.2 Effect of seed-conditioning
We also conducted seed-conditioning experiments as described in §3.1. The seed strings we use are $10$ randomly sampled uppercase characters from the English alphabet. We report the quality-diversity plots in Fig 31 (for next-token prediction on XSUM) and Fig 32 (for multi-token prediction on XSUM). As such, we do not find any changes in diversity, perhaps because this is not a sufficiently open-ended task.
@@ I.2 Effect of seed-conditioning
Figure 31: Seed-conditioning has no effect on diversity for GPT models on XSUM summarization with next-token prediction.
@@
Figure 32: Seed-conditioning has no effect on diversity for GPT models on XSUM summarization with multi-token prediction.
@@
Other studies have proposed benchmarks for evaluating creativity. AidanBench (McLaughlin et al., 2024) and NoveltyBench (Zhang et al., 2025) evaluate LMs on their ability to produce diverse and coherent responses by penalizing repetition across generations. However, they do not measure originality relative to training data, leaving open whether outputs are genuinely novel or simply unseen paraphrases/recombinations. Zhao et al. (2024) evaluate LM creativity using the Torrance Tests of Creative Thinking, a standard in human psychometrics. Another line of work such as Alchemy (Wang et al., 2021), IVRE (Xu et al., 2022), and DiscoveryWorld (Jansen et al., 2024) present simulations with hidden facts and rules, requiring LMs to explore, hypothesize, and test through interaction. While these simulations focus on pretrained models rather than examining how training shapes creative capabilities, they serve as valuable and realistic benchmarks for assessing the role of creativity in scientific discovery.
@@
In support of next-token prediction, there are arguments (Shannon, 1948, 1951; Alabdulmohsin et al., 2024) that claim that language is captured by NTP with models even superceding humans (Shlegeris et al., 2022) at NTP. There are also theoretical results emphasizing the expressivity (Merrill & Sabharwal, 2024; Feng et al., 2023) and learnability (Malach, 2023; Wies et al., 2023) of autoregressive Transformers as long as there is a sufficiently long chain of thought.
@@
While these methods employ diverse strategies, a common feature is their reliance on multi-token objectives that capture broader dependencies across entire sequences. Representative examples include teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) and independent output heads or modules (Gloeckle et al., 2024; DeepSeek-AI et al., 2024) or inserting a lookahead attention (Du et al., 2023). Another line of research is discrete diffusion models (Hoogeboom et al., 2021; Austin et al., 2021; Gong et al., 2023; Lou et al., 2023), which avoid strict left-to-right factorization by iteratively refining an entire sequence at multiple positions. There are other models as well, such as energy-based models (Dawid & LeCun, 2023) and non-autoregressive models or (Gu et al., 2018).
@@
Graph tasks have been used to understand various limitations of Transformers in orthogonal settings. Bachmann & Nagarajan (2024); Saparov et al. (2024) report that Transformers are limited in terms of learning to search tasks on graphs, while Sanford et al. (2024) provide positive expressivity results for a range of algorithmic tasks that process an graph. These works differ from our study of combinational creativity since their graphs are provided in-context and the tasks have a unique answer. Other works (Schnitzler et al., ; Yang et al., 2024a, b) study multi-hop question answering on a knowledge graph; however, this does not require planning.
@@
One line of work relevant to us in the history of generative models is RNN-based VAE for text data (Bowman et al., 2016). The motivation, like in our work, was to learn high-level semantic features rather than next-token features with the hope of producing more novel sentences. However, this suffered from posterior collapse, where the model ignores the latent variable altogether inspiring various solutions (Yang et al., 2017; Goyal et al., 2017). Our results on noise-conditioning are also reminiscent of a line of work on exploration in reinforcement learning (RL), where it has been shown that adding noises to the policy model parameters enables more efficient exploration than directly adding noises to the output space (Plappert et al., 2017; Fortunato et al., 2017).